A body falls freely undergravity from a height of 40m into the top of a platform 0.8m above the ground. It velocity on reaching the platform is_______. Take g=9.8m/s
1
Expert's answer
2019-11-21T10:02:44-0500

The law of conservation of energy gives


"mgh_i=\frac{mv_f^2}{2}+mgh_f"

Thus


"v_f=\sqrt{2g(h_i-h_f)}"

"=\sqrt{2\times 9.8\:\rm m/s^2\times (40\: m-0.8\: m)}=27.7\:\rm m/s"